Given that𝜇=15  and 𝜎=3


Suppose samples of size five are taken. 

1. What is the mean of the sampling distribution?    

2. What is the standard deviation of the sampling distribution?    

(Round answers to TWO decimal places)



1
Expert's answer
2021-07-12T18:48:00-0400

1. Mean="\mu" =15

2. Standard deviation of "\bar X=\frac{\sigma} {\sqrt n}"

"=\frac{3}{\sqrt 5}"

=1.34